Thursday, 20 August

Special Activities

\underline{14:00 – 18:00 H 104}

Mathematical Software Chair: Winfried Neun, ZIB, Berlin, Germany In this session, a broad spectrum of mathematical software including symbolic and numerical systems as well as interfaces and problem solutions will be presented. Details of the presentations can be found on the ICM’98 web server at URL http://elib.zib.de/ICM98/MathSoftware .

14:00 – 14:30 Ewgenij Gawrilow and Michael Joswig, Technische Universität Berlin, Germany Polymake

14:40 – 15:10 Peter Dräxler, Universität Bielefeld, Germany CREP

15:15 – 15:45 Daniel Grayson and Michael Stillman, University of Illinois, USA Macaulay 2

15:45 – 16:15 Break

16:15 – 16:45 Stephen Linton et. al., St Andrews, Scotland and Aachen, Germany GAP

16:50 – 17:20 Walter Oevel, Benno Fuchssteiner et. al., Universität Paderborn, Germany MuPAD

17:30 – 18:00 Fabrice Rouillier et. al., INRIA Lorraine, France FRISCO

ICM’98 – Program  1